# CAIRO: Decoupling Order from Scale in Regression

### v1

@@ Abstract
Standard regression methods typically optimize a single pointwise objective, such as mean squared error, which conflates the learning of ordering with the learning of scale. This coupling renders models vulnerable to outliers and heavy-tailed noise. We propose CAIRO (Calibrate After Initial Rank Ordering), a framework that decouples regression into two distinct stages. In the first stage, we learn a scoring function by minimizing a scale-invariant ranking loss; in the second, we recover the target scale via isotonic regression. We theoretically characterize a class of "Optimal-in-Rank-Order" objectives — including variants of RankNet and Gini covariance — and prove that they recover the ordering of the true conditional mean under mild assumptions. We further show that subsequent monotone calibration guarantees recovery of the true regression function. Empirically, CAIRO combines the representation learning of neural networks with the robustness of rank-based statistics. It matches the performance of state-of-the-art tree ensembles on tabular benchmarks and significantly outperforms standard regression objectives in regimes with heavy-tailed or heteroskedastic noise.
@@ 1 Introduction
In classification settings, this dual requirement is common, for instance, in ad click prediction or medical screening, where models must rank candidates effectively while also outputting calibrated probabilities. However, this need also extends to regression tasks with continuous, real-valued targets. A prime example is insurance pricing: a model must rank policyholders by risk to segment them correctly, but it must also assign precise premiums to ensure profitability to the insurer.
While methodologies for combining calibration and ranking are well-established in classification, they are under-explored for regression with real-valued targets. Consequently, techniques designed for binary or discrete outcomes do not naturally generalize to continuous settings.
In response, we introduce CAIRO (Calibrate After Initial Rank Ordering), a two-stage framework designed to optimize both ranking and regression performance for real-valued targets. CAIRO learns sequentially: Stage 1 minimizes a ranking loss to learn the optimal ordering, and Stage 2 minimizes a pointwise loss to recover the scale. This framework is theoretically grounded; we show that it targets the true conditional expectation at the population level, just as a standard MSE-based learner would.
Additionally, the separation of learning in stages provides a significant advantage over standard regression: because ranking losses are scale-invariant, the representation learned in Stage 1 is inherently shielded from the magnitude of outliers. This allows CAIRO to function as a robust regression method. It is particularly effective in regimes where the data exhibits heavy tails or heteroskedasticity, offering stability where standard MSE minimization often fails.
@@ 2 Background and Related Work
Recalibration of scores.
Many learning algorithms produce scores whose magnitude is not directly interpretable as it is generally not on the same numerical scale as the quantity one ultimately wishes to predict. Calibration refers to the task of learning a transformation of these raw outputs so that they match a desired target scale. The primary objective of probability calibration is to transform the unnormalized outputs of a classifier into reliable probability estimates that reflect the true empirical likelihood of the target event. Among the most established techniques, Platt scaling assumes a parametric relationship, fitting a sigmoid function to map scores to probabilities, which is particularly effective when the distortion is S-shaped (Platt and others, 1999; Niculescu-Mizil and Caruana, 2005). Alternatively, histogram binning takes a non-parametric approach by partitioning the score space into disjoint intervals and assigning the average label frequency within each bin as the calibrated probability, though this introduces discontinuities (Zadrozny and Elkan, 2001; Naeini et al., 2015). Isotonic regression offers a more flexible non-parametric solution by finding the optimal non-decreasing function that minimizes the squared error between the scores and the targets (Zadrozny and Elkan, 2002). From a theoretical standpoint, this can be interpreted as a shape-constrained regression problem, specifically solving for a univariate monotone projection of the data (often via the Pool Adjacent Violators Algorithm), which allows it to correct arbitrary monotonic distortions without assuming a rigid parametric form (Barlow, 1972; Guntuboyina and Sen, 2018).
@@
In this paper, (i) we propose a two-step regression framework that extends the rank-then-calibrate approach of Menon et al. (2012) to real-valued outcomes. To ensure that the two-stage approach recovers the true regression function at the population level, we formalize the requirement that we expect of the first stage ranking loss and group them under a single collection of loss functions. We then convert the resulting, minimizing score into calibrated predictions using one-dimensional isotonic regression. Our approach provides more flexibility than previous works as it accommodates many ranking loss functions. (ii) Furthermore, we provide insight on how to interpret the ranking losses as we connect them directly to popular measures of associations such as Spearman’s $\rho$ and Kendall’s $\tau$. (iii) Finally, we demonstrate that our method may also be interpreted as a robust alternative to MSE minimization which we show through Numerical Studies.
@@ 3 The Decoupled Regression Framework
The purpose of Stage 1 is to learn a monotone transformation of the true regression function $m^{\star}$. Specifically, later, we will show that we expect the optimal score from Stage 1 to be of the form $g^{\star}(X)=h(m^{\star}(X))$ for some $h$ strictly increasing. In Stage 2, the output from Stage 1 is calibrated with isotonic regression to recover $m^{\star}$. Specifically, we apply isotonic regression in Stage 2 which minimizes the squared loss $\mathcal{L}_{\textup{iso}}$ over the space of non-decreasing functions $\mathcal{F}$ while preserving the ranking learned in Stage 1. Formally,
@@ 4 Estimation and Algorithms
Let $\tilde{f}$ denote the resulting solution to (15). To further arrive at a continuous mapping $\hat{f}$, we utilize the implementation from Pedregosa et al. (2011), which linearly interpolates between the original fitted values, the $\hat{f}(\hat{g}(x_{i}))$’s, and clips predictions outside the training range. For prediction we then use
$\hat{m}(x)\equiv\hat{f}\big(\hat{g}(x)\big),$
@@ 6 Discussions
We have introduced CAIRO, a framework that decouples regression into two distinct statistical tasks: learning the correct ordering of predictions and subsequently recovering their scale. By prioritizing the ranking objective, we circumvent the sensitivity of standard pointwise losses, such as Mean Squared Error (MSE), to outliers and heavy-tailed noise. Our theoretical analysis identifies a broad class of Optimal-in-Rank-Order (ORO) loss functions–including smooth approximations of Kendall’s $\tau$ and Gini covariance-that guarantee consistency with the true conditional mean. We further prove that composing these rankers with isotonic calibration recovers the regression function $m^{*}(x)=\mathbb{E}[Y|X=x]$ at the population level. The decoupling of order and scale mainly impacts robustness. As demonstrated in our noise experiments, standard MSE minimization is often driven by a few high-leverage outliers. In contrast, rank-based objectives remain stable because they depend on the sign of pairwise differences or the rank of targets, rather than their absolute magnitude. Our framework suggests a re-evaluation of standard regression practices. In many scientific and industrial applications-such as risk scoring, drug discovery, or prioritization-the relative ordering of candidates is often more critical than their precise numerical values. CAIRO formalizes this intuition, demonstrating that one can often achieve superior regression performance by first solving the ranking problem.

### v2

@@ Abstract
Standard regression methods typically optimize a single pointwise objective, such as mean squared error, which conflates the learning of ordering with the learning of scale. This coupling renders models vulnerable to outliers and heavy-tailed noise. We propose CAIRO (Calibrate After Initial Rank Ordering), a framework that decouples regression into two distinct stages. In the first stage, we learn a scoring function by minimizing a scale-invariant ranking loss; in the second, we recover the target scale via isotonic regression. We theoretically characterize a class of "Optimal-in-Rank-Order" objectives — including variants of RankNet and Gini covariance — and prove that they recover the ordering of the true conditional mean under mild assumptions. We further show that subsequent monotone calibration recovers the true regression function at the population level and mathematically guarantees that finite-sample predictions are strictly auto-calibrated. Empirically, CAIRO combines the representation learning of neural networks with the robustness of rank-based statistics. It matches the performance of state-of-the-art tree ensembles on tabular benchmarks and significantly outperforms standard regression objectives in regimes with heavy-tailed or heteroskedastic noise.
@@ 1 Introduction
In classification settings, this dual requirement is common, for instance, in ad click prediction or medical screening, where models must rank candidates effectively while also outputting calibrated probabilities. However, this need also rigorously extends to regression tasks with continuous, real-valued targets. A prime example is insurance pricing: a model must rank policyholders by risk to segment them correctly, but it must also output strictly auto-calibrated premiums—meaning the expected actual loss for any predicted premium tier must exactly match the predicted premium ($\mathbb{E}[Y|\hat{m}(X)]=\hat{m}(X)$). Without this strict local unbiasedness, an insurer risks severe financial deficits.
While methodologies for combining calibration and ranking are well-established in classification, they are under-explored for regression with real-valued targets. Consequently, techniques designed for binary or discrete outcomes do not naturally generalize to continuous settings. In response, we introduce CAIRO (Calibrate After Initial Rank Ordering), a two-stage framework designed to optimize both ranking and regression performance for real-valued targets. CAIRO learns sequentially: Stage 1 minimizes a ranking loss to learn the optimal ordering, and Stage 2 minimizes a pointwise loss to recover the scale. This framework is theoretically grounded; we show that it targets the true conditional expectation at the population level, just as a standard MSE-based learner would.
Crucially, the separation of learning into distinct stages provides two major statistical advantages over standard unconstrained regression. First, because ranking losses are scale-invariant, the representation learned in Stage 1 is inherently shielded from the magnitude of outliers. This allows CAIRO to function as a highly robust regression method, offering stability in regimes with heavy-tailed or heteroskedastic noise where standard MSE minimization often fails. Second, the isotonic projection in Stage 2 mathematically guarantees that the final predictions are auto-calibrated on the training data. Standard neural networks optimized for global MSE frequently suffer from local miscalibration, systematically over- or under-predicting across different segments of the output space. By structurally enforcing that the empirical mean of the targets exactly matches the predicted value within every constant piece (level set) of the fitted function, CAIRO ensures locally unbiased estimation.
@@ 2 Background and Related Work
Recalibration and auto-calibration.
Many learning algorithms produce scores whose magnitude is not directly interpretable as it is generally not on the same numerical scale as the quantity one ultimately wishes to predict. Calibration refers to the task of learning a transformation of these raw outputs so that they match a desired target scale. In classification, the primary objective is to transform unnormalized outputs into reliable probability estimates via parametric mappings like Platt scaling (Platt and others, 1999; Niculescu-Mizil and Caruana, 2005) or non-parametric approaches like histogram binning (Zadrozny and Elkan, 2001; Naeini et al., 2015). In continuous regression, the analogous concept is auto-calibration (Krüger and Ziegel, 2021; Gneiting and Resin, 2023; Wüthrich, 2023), which requires the prediction to be unbiased conditional on itself, i.e., $\mathbb{E}[Y|\hat{m}(X)]=\hat{m}(X)$. Isotonic regression offers a flexible non-parametric solution to achieve this by finding the optimal non-decreasing function that minimizes the squared error between scores and targets (Zadrozny and Elkan, 2002). Because the Pool Adjacent Violators (PAV) algorithm computes block-wise local averages of the response, isotonic regression structurally guarantees empirical auto-calibration without assuming a rigid parametric form (Barlow, 1972; Guntuboyina and Sen, 2018).
@@
Semiparametric single-index models.
From a statistical perspective, modeling the conditional mean as $\mathbb{E}[Y|X=x]=f(g(x))$, where $f$ is an unknown monotonic link function and $g$ is a scoring index, naturally connects our framework to semiparametric single-index models (SIMs) (Härdle et al., 1993; Härdle and Stoker, 1989; Ichimura, 1993). Our approach shares conceptual roots with the maximum rank correlation (MRC) estimator (Han, 1987), which recovers a linear index $g(x)=x^{\top}\beta$ by maximizing Kendall’s $\tau$. CAIRO thus serves as a flexible, deep-learning-compatible generalization of classical index models. However, it significantly departs from standard SIM estimation strategies. While traditional SIMs rely on computationally heavy profile-likelihoods or bandwidth-sensitive kernel smoothing to jointly estimate $f$ and $g$, CAIRO decouples the problem. By learning a highly non-linear index via distribution-free ranking losses and subsequently recovering the shape-constrained link through efficient isotonic regression, we circumvent classical computational bottlenecks.
@@
In this paper, (i) we propose CAIRO, a two-step regression framework that extends the rank-then-calibrate approach of Menon et al. (2012) to real-valued outcomes. To ensure that the two-stage approach recovers the true regression function at the population level, we formalize the requirement that we expect of the first stage ranking loss and group them under a single collection of Optimal-in-Rank-Order (ORO) loss functions. (ii) We provide theoretical insights into these ranking losses by explicitly connecting them to popular measures of association such as Spearman’s $\rho$ and Kendall’s $\tau$. (iii) We mathematically ensure that our final estimator is auto-calibrated via Stage 2 isotonic projection, effectively turning uncalibrated neural network representations into reliable, locally unbiased statistical forecasters. (iv) Finally, we demonstrate empirically that the decoupling of scale and order allows CAIRO to function as a highly robust alternative to MSE minimization, exhibiting significant resilience to heavy-tailed and heteroskedastic noise.
@@ 3 The Decoupled Regression Framework
The purpose of Stage 1 is to learn a strictly monotone transformation of the true regression function $m^{\star}$. Specifically, later, we will show that we expect the optimal score from Stage 1 to be of the form $g^{\star}(X)=h(m^{\star}(X))$ for some $h$ strictly increasing. In Stage 2, the output from Stage 1 is calibrated with isotonic regression to recover $m^{\star}$. Specifically, we apply isotonic regression in Stage 2 which minimizes the squared loss $\mathcal{L}_{\textup{iso}}$ over the space of non-decreasing functions $\mathcal{F}$ while preserving the ranking learned in Stage 1. Formally,
@@ 3.2 Stage 2: Calibration via Isotonic Regression
Remark 2 (Connection to auto-calibration and proper scoring).
A prediction function $m(X)$ is strictly auto-calibrated if it satisfies $m(X)=\mathbb{E}[Y|m(X)]$ (Krüger and Ziegel, 2021; Gneiting and Resin, 2023). The true conditional mean $m^{\star}(X)=\mathbb{E}[Y|X]$ naturally satisfies this property: $\mathbb{E}[Y|m^{\star}(X)]=\mathbb{E}[\mathbb{E}[Y|X]\mid m^{\star}(X)]=m^{\star}(X).$ While the Stage 1 ranking objectives (such as the Gini covariance) excel at learning robust, scale-invariant representations, they are fundamentally improper scoring rules when evaluated over unconstrained function classes; their raw outputs $g^{\star}(X)$ do not correspond to the true conditional expectation and are therefore not auto-calibrated. Stage 2 resolves this limitation by solving the monotone least-squares problem, isotonic regression explicitly estimates the conditional expectation $\mathbb{E}[Y|g^{\star}(X)]$ while preserving the order learned in Stage 1. As established in Theorem 3, composing the uncalibrated ranker with this monotonic projection recovers $m^{\star}(X)$.
Consequently, CAIRO mathematically guarantees that the final predictions are auto-calibrated, effectively converting an improper ranking mechanism into a statistically rigorous regression estimator. Crucially, because isotonic regression estimates conditional expectations via local block-averaging, this auto-calibration property holds not just asymptotically, but strictly on the empirical training data, even when the Stage 1 ranker is estimated with finite-sample noise.
@@ 4 Estimation and Algorithms
Let $\tilde{f}$ denote the resulting solution to (15). Because the Pool Adjacent Violators (PAV) algorithm fits piecewise-constant steps by taking the exact empirical average of the responses $y_{i}$ within each block, the resulting predictor $\hat{m}(x_{i})=\tilde{f}(\hat{g}(x_{i}))$ is guaranteed to be strictly empirically auto-calibrated on the training set. This provides a robust statistical safety net: even if the finite-sample estimate $\hat{g}$ deviates from the true optimal ordering, the final predictions remain locally unbiased with respect to the learned representation. To further arrive at a continuous mapping $\hat{f}$ for out-of-sample prediction, we utilize the implementation from Pedregosa et al. (2011), which linearly interpolates between the original fitted values, the $\tilde{f}(\hat{g}(x_{i}))$’s, and clips predictions outside the training range.
@@ 6 Discussions
We have introduced CAIRO, a framework that decouples regression into two distinct statistical tasks: learning the correct ordering of predictions and subsequently recovering their scale. By prioritizing the ranking objective, we circumvent the sensitivity of standard pointwise losses, such as Mean Squared Error (MSE), to outliers and heavy-tailed noise. Our theoretical analysis identifies a broad class of Optimal-in-Rank-Order (ORO) loss functions—including smooth approximations of Kendall’s $\tau$ and Gini covariance—that guarantee consistency with the true conditional mean. We further prove that composing these rankers with isotonic calibration recovers the regression function $m^{*}(x)=\mathbb{E}[Y|X=x]$ at the population level.
The decoupling of order and scale provides two statistical benefits: robustness and auto-calibration. First, as demonstrated in our noise experiments, standard MSE minimization is often driven by a few high-leverage outliers. In contrast, rank-based objectives remain stable because they depend on the sign of pairwise differences or the rank of targets, rather than absolute magnitudes. Second, standard deep neural networks trained via MSE frequently suffer from local miscalibration. By deferring scale recovery to a Stage 2 isotonic projection, CAIRO structurally forces the predictions to be strictly empirically auto-calibrated. This converts flexible but historically uncalibrated neural representations into reliable, locally unbiased statistical forecasters. Our framework suggests a re-evaluation of standard regression practices.